A mail-order house uses 17,650 boxes a year. Carrying costs are 60 cents per box a year, and ordering costs are $96. The following price schedule applies.


Number of Boxes Price per Box
1,000 to 1,999 $1.25                
2,000 to 4,999 1.20                
5,000 to 9,999 1.15                
10,000 or more 1.10             
Answer #1

DEMAND = 17650

ORDERING COST = 96

HOLDING COST = 0.6

THEORY

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST)

Q* = ADJUSTED QUANTITIY BETWEEN THE UPPER AND LOWER LIMIT AND THE EOQ

AHC = ANNUAL HOLDING COST = (Q* / 2) * HOLDING COST PER UNIT

AOC = ANNUAL ORDERING COST = (DEMAND / Q*) * ORDERING COST

APC = ANNUAL PURCHASING COST = DEMAND * Q*

TCI = TOTAL COST OF INVENTORY = AHC + AOC + APC


NO.

LOWER LIMIT

UPPER LIMIT

PER UNIT

HOLDING COST

EOQ

Q*

AHC

AOC

APC

TCI

1

0

1999

1.25

0.6

2377

1999

(1999 / 2) * 0.6 = 599.7

(17650 / 1999) * 96 = 847.62

17650 * 1.25 = 22062.5

599.7 + 847.62 + 22062.5 = 23509.82

2

2000

4999

1.2

0.6

2377

2377

(2377 / 2) * 0.6 = 713.1

(17650 / 2377) * 96 = 712.83

17650 * 1.2 = 21180

713.1 + 712.83 + 21180 = 22605.93

3

5000

9999

1.15

0.6

2377

5000

(5000 / 2) * 0.6 = 1500

(17650 / 5000) * 96 = 338.88

17650 * 1.15 = 20297.5

1500 + 338.88 + 20297.5 = 22136.38

4

10000

& more

1.1

0.6

2377

10000

(10000 / 2) * 0.6 = 3000

(17650 / 10000) * 96 = 169.44

17650 * 1.1 = 19415

3000 + 169.44 + 19415 = 22584.44


OPTIMAL ORDER QUANTITY = 5000

TOTAL COST OF INVENTORY = 22136